OpenStudy (anonymous):

2+5x = 11 + 8x+3 ---- -- ---- 4 28 7

OpenStudy (anonymous):

x= 3

OpenStudy (anonymous):

would you be able to kind of work out the problem for me? I would appreciate lots

OpenStudy (anonymous):

u wanted to find x right, then sure

OpenStudy (anonymous):

samsterz, do not give direct answers. The lowest common denominator is 28 so multiply everything by that. This would give you 7(2+5x) = 11 + 4(8x+3). From there you can distribute and get 14 + 35x = 11 +32x + 12 which would become 14+35x = 23 + 32x. Get variable on one side by manipulating with subtraction. You will get 14 + 3x = 23. Now get numbers without variables to other side. You will get 3x = 9. Get x alone, x=3.

OpenStudy (anonymous):

ash vasan i said id give an explanation

OpenStudy (anonymous):

ok back

OpenStudy (anonymous):

Multiply each term by a factor of 1 that will equate all the denominators. In this case, all terms need a denominator of 28. The (8x+3)7 expression needs to be multiplied by (4)(4) to make the denominator 28. 1/4(2+5x)=1128+17(8x+3)⋅44 Multiply the expression by a factor of 1 to create the least common denominator (LCD) of 28. 1/4(2+5x)=1128+(8x+3)(4)28 Multiply 4 by each term in 8x+3 to get 32x+12. Multiply each term in the first polynomial by each term in the second polynomial. 1/4(2+5x)=1128+8x⋅4+3⋅428 Multiply 8x by 4 to get 32x. 1/4(2+5x)=1128+128(32x+3⋅4) Multiply 3 by 4 to get 12. 1/4(2+5x)=1128+128(32x+12) 1/4(2+5x)=1128+128(32x+12) The numerators of expressions that have equal denominators can be combined. In this case, 1128 and (32x+12)28 have the same denominator of 28, so the numerators can be combined. 1/4(2+5x)=11+(32x+12)28 Simplify the numerator of the expression. 1/4(2+5x)=128(11+32x+12) Add 12 to 11 to get 23. 1/4(2+5x)=128(23+32x) Reorder the polynomial 23+32x alphabetically from left to right, starting with the highest order term. 1/4(2+5x)=128(32x+23) Reorder the polynomial 2+5x alphabetically from left to right, starting with the highest order term. 1/4(5x+2)=128(32x+23) Since there is one rational expression on each side of the equation, this can be solved as a ratio. For example, AB=CD is equivalent to A⋅D=B⋅C. (5x+2)⋅28=(32x+23)⋅4 Multiply 4 by each term inside the parentheses. More Detail Multiply (32x+23) by 4 to get 4(32x+23). (5x+2)⋅28=4(32x+23) Multiply 4 by each term inside the parentheses (32x+23). (5x+2)⋅28=4(32x)+4(23) Multiply 4 by the 32x inside the parentheses. (5x+2)⋅28=4⋅32x+4(23) Multiply 4 by 32x to get 128x. (5x+2)⋅28=128x+4(23) Multiply 4 by the 23 inside the parentheses. (5x+2)⋅28=128x+4⋅23 Multiply 4 by 23 to get 92. (5x+2)⋅28=128x+92 (5x+2)⋅28=128x+92 Multiply 28 by each term inside the parentheses. Multiply (5x+2) by 28 to get 28(5x+2). 28(5x+2)=128x+92 Multiply 28 by each term inside the parentheses (5x+2). 28(5x)+28(2)=128x+92 Multiply 28 by the 5x inside the parentheses. 28⋅5x+28(2)=128x+92 Multiply 28 by 5x to get 140x. 140x+28(2)=128x+92 Multiply 28 by the 2 inside the parentheses. 140x+28⋅2=128x+92 Multiply 28 by 2 to get 56. 140x+56=128x+92 140x+56=128x+92 Since 128x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 128x from both sides. 140x+56−128x=92 Since 140x and −128x are like terms, add −128x to 140x to get 12x. 12x+56=92 Since 56 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 56 from both sides. 12x=−56+92 Add 92 to −56 to get 36. 12x=36 Divide each term in the equation by 12. Simplify the left-hand side of the equation by canceling the common factors. Cancel the common factor of 12 in 12x/12. Reduce the expression 12x/12 by removing a factor of 12 from the numerator and denominator. x=36/12 Simplify the equation. Cancel the common factor of 12 in 36/1212. Reduce the expression 3612 by removing a factor of 12 from the numerator and denominator.

OpenStudy (anonymous):

and u get x=3

OpenStudy (anonymous):

a lot of thanks to both